OpenStudy (anonymous):

How to find the derivative of cos^2(9x)

OpenStudy (anonymous):

chain rule. hint: cos^2(9x) = [ cos(9x) ]^2

OpenStudy (anonymous):

I'm still confused, can I have a little more help please? haha

OpenStudy (anonymous):

For the chain rule, let f(x) = cos^2(x) and g(x) = 9x Then f'(x) = f'(cos(x)*cos(x)) = 2sin(x)cos(x) and g'(x) = 9 So you get f'(g(x))*g'(x) = 2sin(9x)cos(9x)*9 = 18sin(9x)cos(9x)